Operations Report – January 31st

Crew 240 Operations Report 31-01-2022

SOL: 0

Name of person filing report: François Vinet

Non-nominal systems: Heater

Notes on non-nominal systems: The temperature has dropped a lot from the night in the upper desk (3am : 61°F, 6am : 57°F). Atila checked the thermostat and switched to FAN:AUTO (it wasn’t the case during the night). Filter was also replaced.


Spirit rover used: yes

Hours: 153.7

Beginning charge: 100%

Ending charge: 76%

Currently charging: yes

Opportunity rover used: yes

Hours: 82.9

Beginning charge: 100%

Ending charge: 70%

Currently charging: yes

Curiosity rover used: yes

Hours: 165.4

Beginning charge: 100%

Ending charge: 70%

Currently charging: yes

Perseverance rover used: No

Hours: 226.5

Beginning charge: NA

Ending charge: 100%

Currently charging: NA

General notes and comments: NA

Summary of Hab operations: Some crew members felt discomfort during the night because of dusty air. The air filter has been changed. Water filter next to the loft tank has also been changed.

WATER USE: 11 gallons

Water (static tank): 470 gallons

Water (loft tank): 24 gallons

Water Meter: 0155318.2 units

Static to Loft Pump used : yes

Static tank pipe heater (on or off): on

Static tank heater (On or off): on

Toilet tank emptied: no

Summary of internet: NA

Summary of suits and radios: NA

Summary of GreenHab operations:

WATER USE: 10 gallons

Heater: Off

Supplemental light: Off

Harvest: Cherry tomatoes (150g)


– Should the supplemental light be turned on during our weeks at the MDRS considering the sun sets early?

– If the GreenHab is at 80°F at 7PM should the heater be turned on considering the temperature will likely drop even lower overnight?

Summary of ScienceDome operations:

Dual split: Off

SOC : 41% at 6.45pm

Summary of RAM operations: The batteries of the Combustible Gas Meter have been changed (3 AAA batteries).

Summary of any observatory issues: The computer in the observatory is not working (problem at startup). Peter Detterline has been contacted by the crew astronomer on his personal gmail address. The main computer mouse is also broken (there is no left key anymore).

Summary of health and safety issues: There is a problem in the insulation of the commander bedroom: the glass wool above the headboard is apparent. Photos can be provided.

Questions, concerns and requests to Mission Support: one non-burnable trash is waiting for pickup in the rear airlock. Thank you!

– I am not sure what is the dual split in the Science Dome. Is it the white fan to the right of the battery management device ? Where can I check its state ?

Copyright © The Mars Society. All rights reserved. | Main Site